Grade 12Physical Chemistry

Calculate energy of one mole of photons of radiation whose frequency is 5 ×1014 Hz

sudhanshu
12 Years agoGrade 12

2 Answers

Gaurav
11 Years ago
Hello Students
Energy of one photon is given by the expression E = hV
where, h = PlancKs constant = 6.636 x 10-34 J s-1
V= Frequency of radiation = 5x 1014 Hz= 5 x 1014 s-1.
E= (6.636 x 1034 J s-1 x 5 x 1014 s-1)= 33.18 x 1020J
One mole of photon = 6.023 x 1023 photons
Therefore, Energy of one mole of photon = (6.023 x 1023 x 33.18 x 10-20 J) = 199.84 x 103 J
Energy of one mole of photon = 199.84 kJ
Yash Chourasiya
6 Years ago
Hello Students

The energy of one photon (E)=hv

Here,h = 6.626×10-34Js

v = 5×1014s−1

E = 6.626×10−34× 5×1014
= 3.313×10−19J
Thus the energy of one mole of photons isE×A
Where A is the Avogadro number
= 3.313×10−19J× 6.022×1023mol-1

= 199.51 kJmol-1

I hope this answer will help you.